OpenStudy (anonymous):

(x_0)^3+(y_0)^3=(z_0)^3. if 2p=x_0+y_0 and 2q=x_0-y_0, x_0=? and y_0=?

OpenStudy (anonymous):

Adding \(x_0+y_0=2p \text{ and } x_0-y_0=2q\) gives \(2x_0=2p+2q \implies x_0=p+q\). Substitute that in either equation to get \(y_0=2p-x_0=2p-p-q \implies y_0=p-q.\)

OpenStudy (anonymous):

THANK YOU :)

OpenStudy (anonymous):

You're welcome!